OpenStudy (anonymous):

Solve by taking the a logarithm of both sides. 8x = 23 x = 1.21 x = 1.51 x = 1 x = 2 x = 1.21 x = 1.51 x = 1 x = 2

OpenStudy (anonymous):

what are all those things below, choices?

OpenStudy (anonymous):

\[b^x=A\iff x=\frac{\ln(A)}{\ln(b)}\]

OpenStudy (anonymous):

sorry x = 1.21 x = 1.51 x = 1 x = 2 those are the choices

OpenStudy (anonymous):

i take it the equation is \[8^x=23\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so, clearly the last two options are out, since \(8^1=8\) and \(8^2=64\)

OpenStudy (anonymous):

take a calculator and compute \[\frac{\ln(23)}{\ln(8)}\] for your answer

OpenStudy (anonymous):

it wont let me